OpenStudy (anonymous):

15x + 24 = 8(10 + 3x) - 2

OpenStudy (anonymous):

15x +24 = 80 + 24x - 2 15x - 24x = 80 -2 -24 -9x = 54 x = -6

OpenStudy (anonymous):

15x + 24 = 8(10 + 3x) - 2 15x + 24 = 80 + 3x - 2 -3x -3x 12x + 24 = 80 - 2 12x + 24 = 78 -24 -24 12x = 54 Divide both sides by 12 x = 4.5

OpenStudy (anonymous):

@Fang02 in incorrect

OpenStudy (anonymous):

What did I do wrong?

OpenStudy (anonymous):

@coldground93 is correct. @Fang02 you need to distribute the 8 to the 3x as well.

OpenStudy (anonymous):

Oh okay, I forgot about that. Thanks for letting me know!